OpenStudy (anonymous):

Solve. x2 + 10x = –2

OpenStudy (anonymous):

x^2 + 10x + 2 = 0 Use the quadratic formula a = 1 b = 10 c = 2 x = -5-sqrt(23) x = sqrt(23)-5

OpenStudy (anonymous):

\[x^2+10x+2=0\] use quadratic formula \[(-b \pm \sqrt{b^2-4ac})2a\] where a=1,b=10 and c=2

OpenStudy (anonymous):

sorry, its divided by 2a